The diameter of the rear and front wheels of a pushcart are 30 and 18 respectively. In what distance in feet would the front wheel make 20 revolutions more than rear wheels?
Answer by ankor@dixie-net.com(22740)

The diameter of the rear and front wheels of a pushcart are 30 and 18 respectively.
In what distance in feet would the front wheel make 20 revolutions more than rear wheels?
:
The ratio of the two circumferences is the same as the two diameters: 30:18
let n = no. of rev of the rear wheel
30n = 18(n+20)
30n = 18n + 360
30n - 18n = 360
n = 360/12
n = 30 rev for the rear wheel, to exceed the front wheel rev by 20
:
Find the distance using the circumference
30*pi*30 = 2827.4 inches, back wheel
:
Check this using the front wheel
the front wheel goes the same dist when it has 50 rev
50*pi*18 = 2827.4 inches
Find this distance in ft
2827.4/12 = 235.6 ft